# Put CASH on Bandits: A Max K-Armed Problem for Automated Machine Learning

### v1

@@ 5 Performance on AutoML tasks
Next, we examine the empirical performance of MaxUCB in an AutoML setting via reporting average ranking and the number of wins, ties, and losses across tasks for each benchmark (details in Appendix D.1). We first focus on the impact of the hyperparameter $\alpha$ and then compare our approach to others. Specifically, we show that our two-level approach performs better than single-level HPO on the joint space, and MaxUCB outperforms other state-of-the-art bandit methods. To begin, we will provide a brief overview of the experimental setup used across all experiments.
Experimental setup. We use four AutoML benchmarks [^4] We used the AutoML Toolkit (AMLTK) (Bergman et al., 2024). We ran HPO on a compute cluster equipped with Intel Xeon Gold 6 24062406\,2406 240 CPUs, requiring 20 0002000020\,00020 000 CPU hours. We conducted the remaining experiments on a local machine with Intel Core i7-1370P, requiring an additional $32$ CPU hours. , all implementing CASH for tabular supervised learning differing in the considered ML models, HPO method, and datasets, which we detail in Table D.1.
For TabRepo and Reshuffling, we use available pre-computed HPO trajectories, whereas for TabRepoRaw and YaHPOGym, we use SMAC (Hutter et al., 2011; Lindauer et al., 2022) as Bayesian optimization method to run HPO ourselves.
@@
How sensitive is MaxUCB to the choice of $\alpha$? We study sensitivity to $\alpha$ in Figure 5, showing that the choice of $\alpha$ impacts performance. Lower values of $\alpha$ (light red lines) lead to good performance with small budgets, whereas higher values (dark red lines) achieve stronger final performance when sufficient time is available. An $\alpha$ around $0.5$ yields a balanced trade-off, offering robust anytime performance across tasks (see Appendix D.4 for detailed results). This analysis also indicates the possibility of improving performance by meta-learning how to set $\alpha$ for a new dataset based on its characteristics and the available time budget. In particular, $\alpha=0.5$ is recommended when the reward distribution support is $[0,1]$. Competitive Baselines. We compare against Quantile Bayes UCB (Balef et al., 2024), ER-UCB-S (Hu et al., 2021) and Rising Bandits (Li et al., 2020) which have been developed for the decomposed CASH task. We consider extreme bandits (QoMax-SDA (Baudry et al., 2022), Max-Median (Bhatt et al., 2022)) and classic UCB as general bandit methods. We use default hyperparameter settings for all methods. As combined search baselines, we consider Bayesian optimization (SMAC) and random search.555 Only available for TabRepoRaw and YaHPOGym, where we computed HPO trajectories ourselves. Additionally, we report the performance of the best (oracle) arm.
How does the two-level approach compare against combined search?
@@
We first compare the average rank over time in Figure 6 using combined search (black lines; random search and SMAC if available) to the two-level approach. We observe that all methods outperform random search (dotted line) and that while SMAC quickly catches up, some bandit methods continuously achieve a better (lower) rank. Additionally, we observe that any of the mentioned MAB algorithms (except Rising Bandits and QoMax-SDA) leads to superior performance in the early stages ($T=50$). This demonstrates that the decomposition is particularly useful when the number of iterations is limited. Additionally, Table 1 shows that the difference in final performance between combined search and the two-level approach is significant.
How does MaxUCB compare against other bandit methods? Figure 6 shows a substantial difference in the ranking.
@@
In the beginning, many methods perform competitively, but MaxUCB yields the best anytime and final performance. Concretely, classical UCB (red) and extreme bandits (QoMax-SDA; brown) perform worst. The Max-Median algorithm (purple) shows strong initial performance, but its effectiveness declines with more iterations. While Max-Median identifies and avoids the worst arm; it sometimes struggles to select the optimal arm, resulting in non-robust performance.
Next, we look at methods that were originally designed for AutoML. Both ER-UCB-S (pink) and Quantile Bayes UCB (orange) focus on estimating the higher region of the reward distribution rather than the extreme values. ER-UCB-S assumes a Gaussian reward distribution and is consistently outperformed by Quantile Bayes UCB, a distribution-free algorithm. Rising Bandits (green) underperforms initially (requiring $KC$ steps, where $K$ is #arms, and $C$ is a hyperparameter set to $7$) but reaches a competitive final performance.
@@
This is especially pronounced for the YaHPOGym benchmark, where Rising Bandits outperforms MaxUCB with respect to normalized average performance (see Figure D.5 in Appendix D.5). This benchmark contains datasets where the optimal arm changes over time. Since Rising Bandits considers non-stationarity, it performs better for these instances. [^6] A burn-in phase, i.e., pulling each arm for a few rounds at the beginning without observing the rewards, yields a competitive solution as we assess in Appendix E.3.
Finally, MaxUCB and Quantile Bayes UCB are the only ones that significantly outperform combined search in Table 1. And looking at TabRepoRaw and Reshuffling, as depicted in Figure 6, demonstrates that MaxUCB is a robust method for CASH problems.
@@
6 Discussion and Future Work
This paper addresses the CASH problem, proposing MaxUCB, an MKB method. Our data-driven analysis answers an open question on the applicability of extreme bandits to CASH. We provide a novel theoretical analysis and show state-of-the-art performance on several important benchmarks.
Limitations. Though our method can be applied beyond AutoML in principle, it is finely tuned for this setting with bounded and skewed distributions and for maximal value optimization. Our analysis relies on stationary distributions, which might not always be accurate, especially at the beginning of the HPO run, so a short burn-in phase may be needed to reach this regime. Lastly, we provide a default value for our hyperparameter $\alpha$ that might need adjusting for other applications.
Impact on AutoML systems. Our approach complements prior work on AutoML systems and increases their flexibility. Firstly, using a combined search with a single HPO process for CASH may restrict the choice of HPO methods. Thus, a decomposed approach allows to easily integrate recent progress in HPO methods for some ML model classes, e.g., multi-fidelity or meta-learned methods  (Li et al., 2018; Falkner et al., 2018; Müller et al., 2023; Chen et al., 2022). Secondly, some AutoML systems (Swearingen et al., 2017; Li et al., 2023) decompose the search space into smaller subspaces to scale to a distributed setting and use Bandit methods to select promising subspaces (Levine et al., 2017; Li et al., 2020). While we focus on applying bandits to select promising ML models, our methods could also be applied in this setting.
Moreover, MaxUCB’s application extends beyond the CASH problem. For example, MaxUCB is well-suited for sub-supernet selection in Neural Architecture Search (NAS) (Hu et al., 2022), where we observe similar reward distributions (Ly-Manson et al., 2024) (see Appendix E.5).
@@ 6 Discussion and Future Work
Future Directions. Future research could explore leveraging trends in reward sequences, which may reveal valuable patterns, especially in non-stationary settings like AutoML for data streams, where optimal models shift with data distributions (Verma et al., 2024). Incorporating cost-aware optimization is another promising direction, as computational resources and time, rather than iteration counts, often define budgets in AutoML; this would require estimating model training times and factoring them into the decision process. Addressing the growing complexity of heterogeneous ML tasks, such as those involving pre-training, fine-tuning, or prompt engineering, may benefit from a hierarchical approach that allocates resources effectively across diverse pipelines. Additionally, exploiting structural similarities among algorithms and their hyperparameters could reduce exploration costs by sharing information across arms in the CASH problem, further enhancing efficiency.
@@ Appendix A Preliminiaries
Let $F(x)=P(X\leq x)$ be the cumulative distribution function, so $G(x)=1-F(x)=P(X>x)$. First, consider the probability that the maximum of the $n$ samples is less than or equal to $x$:
@@ Appendix B Proofs
By applying the Mean Value Theorem (MVT) to the survival function $G$ over an interval $[b-\epsilon,b]$, there exists a point $c\in(b-\epsilon,b)$ such that:
@@
Where $N_{i}(T)=\mathbbm{E}(\sum\limits_{t=1}^{T}\mathbbm{1}{\{I_{t}=i\}})$ is the number of sub-optimal pulls of arm $i$, arm $i^{*}$ is the optimal arm and $b_{i}$ is the upper bound on the reward of arm $i$ as given by Assumption 3.2, i.e., the reward of arm $i$ lies within the interval $[a_{i},b_{i}]$.
@@
=$=\left\{\max_{1\leq n\leq n_{1}(t)}r_{1,n}\leq\mathop{\mathbbm{E}[\max_{t\leq T}r_{1,t}]}-\Delta_{i}-C_{t}(n_{1}(t))\right\}$
@@
  • Underestimating the upper confidence bound of extreme values for the optimal arm $1$
@@
Now we use Lemma A.1 to calculate the probability of 29:
@@
This selection of the function of extrapolation bonus results in two significant advantages. First, it provides an upper bound for the right-hand side of equation 31 that remains independent of $n_{1}$. Furthermore, equation 32 shows a decreasing trend as the number of pulls for the sub-optimal arm $i$ increases.

### v2

@@ 5 Performance on AutoML tasks
Finally, we examine the empirical performance of MaxUCB in an AutoML setting via reporting average ranking and the number of wins, ties, and losses across tasks for each benchmark (details in Appendix D.1). We first focus on the impact of the hyperparameter $\alpha$ and then compare our approach to others. Specifically, we show that our two-level approach performs better than single-level HPO on the joint space, and MaxUCB outperforms other state-of-the-art bandit methods. To begin, we will provide a brief overview of the experimental setup used across all experiments.
Experimental setup. We use four AutoML benchmarks [^5] We used the AutoML Toolkit (AMLTK) (Bergman et al., 2024). We ran HPO on a compute cluster equipped with Intel Xeon Gold 6 2406\,240 CPUs, requiring 20 00020\,000 CPU hours. We conducted the remaining experiments on a local machine with Intel Core i7-1370P, requiring an additional $32$ CPU hours. , all implementing CASH for tabular supervised learning differing in the considered ML models, HPO method, and datasets, which we detail in Table D.1. For TabRepo and Reshuffling, we use available pre-computed HPO trajectories, whereas for TabRepoRaw and YaHPOGym, we use SMAC (Hutter et al., 2011; Lindauer et al., 2022) as a Bayesian optimization method to run HPO ourselves.
Competitive Baselines. We compare against Quantile Bayes UCB (Balef et al., 2024), ER-UCB-S (Hu et al., 2021) and Rising Bandits (Li et al., 2020) which have been developed for the decomposed CASH task. We consider extreme bandits (QoMax-SDA (Baudry et al., 2022), Max-Median (Bhatt et al., 2022)) and classic UCB as general bandit methods. We use default hyperparameter settings for all methods. As combined search baselines, we consider Bayesian optimization (SMAC) and random search. [^6] Only available for TabRepoRaw and YaHPOGym, where we computed HPO trajectories ourselves. Additionally, we report the performance of the best (oracle) arm.
@@
How sensitive is MaxUCB to the choice of $\alpha$? Figure 5, shows that the choice of $\alpha$ impacts performance. Lower values of $\alpha$ (light red lines) lead to good performance with small budgets, whereas higher values (dark red lines) achieve stronger final performance when sufficient time is available. An $\alpha$ around $0.5$ yields a balanced trade-off, offering robust anytime performance across tasks (see Appendix D.4 for detailed results). Another insight from this study is that the right choice of $\alpha$ may depend on characteristics of the datasets, such as the support of the losses, and could be meta-learned.
@@
How does the two-level approach compare against combined search? We first compare the average rank over time in Figure 6 using combined search (black lines; random search and SMAC if available) to the two-level approach. We observe that all methods outperform random search (dotted line) and that while SMAC quickly catches up, some bandit methods continuously achieve a better (lower) rank. Additionally, we observe that most MAB algorithms (except Rising Bandits and QoMax-SDA) lead to superior performance in the early stages ($T=50$). This demonstrates that the decomposition is particularly useful when the number of iterations is limited. Additionally, Table 1 shows that the difference in final performance between combined search and the two-level approach is significant.
@@
How does MaxUCB compare against other bandit methods? Figure 6 shows a substantial difference in the ranking. In the beginning, many methods perform competitively, but MaxUCB yields the best anytime and final performance. Classical UCB (red) and extreme bandits (QoMax-SDA; brown) perform worst. The Max-Median algorithm (purple) shows strong initial performance, but its effectiveness declines with more iterations. While Max-Median identifies and avoids the worst arm; it sometimes struggles to select the optimal arm, resulting in non-robust performance.
@@
Next, we look at methods that were originally designed for AutoML. Both ER-UCB-S (pink) and Quantile Bayes UCB (orange) focus on estimating the higher region of the reward distribution rather than the extreme values. ER-UCB-S assumes a Gaussian reward distribution and is consistently outperformed by Quantile Bayes UCB, a distribution-free algorithm. Rising Bandits (green) underperforms initially due to its costly initialization but reaches a competitive final performance. This is especially pronounced for the YaHPOGym benchmark, where Rising Bandits outperforms MaxUCB with respect to normalized average performance (see Figure D.5 in Appendix D.5). This benchmark contains datasets where the optimal arm changes over time. Since Rising Bandits models non-stationary rewards, it performs better for these instances. [^7] A burn-in phase, i.e., pulling each arm for a few rounds at the beginning without observing the rewards, yields a competitive solution as we assess in Appendix E.3. Finally, MaxUCB and Quantile Bayes UCB are the only ones that significantly outperform combined search in Table 1. And looking at TabRepoRaw and Reshuffling, as depicted in Figure 6, demonstrates that MaxUCB is a robust method for CASH problems.
@@
6 Conclusions, Discussions and Future Work
This paper addresses the CASH problem, proposing MaxUCB, an MKB method. Our data-driven analysis answers an open question on the applicability of extreme bandits to CASH. We provide a novel theoretical analysis and show state-of-the-art performance on several important benchmarks. Limitations. Though our method can be applied beyond AutoML in principle, it is finely tuned for this setting with bounded and skewed distributions and for maximal value optimization. Our analysis relies on stationary distributions, which might not always be accurate, especially at the beginning of the HPO run, so a short burn-in phase may be needed to reach this regime. Our approach may not be distributionally optimal, as optimality in bounded extreme bandits remains an open question, and establishing lower bounds is left for future work. Lastly, we provide a default value for our hyperparameter $\alpha$ that might need adjusting for other applications.
Impact on AutoML systems. Our approach complements prior work on AutoML systems and increases their flexibility. First, our approach allows to choose any HPO method for each model at the lower level and thus it may integrate recent progress in HPO methods for some ML model classes, e.g., multi-fidelity or meta-learned methods  (György and Kocsis, 2011; Li et al., 2018; Falkner et al., 2018; Müller et al., 2023; Chen et al., 2022). Second, some AutoML systems (Swearingen et al., 2017; Li et al., 2023) decompose the search space into smaller subspaces to scale to a distributed setting and use Bandit methods to select promising subspaces (Levine et al., 2017; Li et al., 2020). While we focus on applying bandits to select promising ML models, our methods could also be applied in this setting. Finally, beyond CASH, MaxUCB is well-suited for sub-supernet selection in Neural Architecture Search (NAS) (Hu et al., 2022), showing similar reward distributions (Ly-Manson et al., 2024) (see Appendix E.5).
@@ 6 Conclusions, Discussions and Future Work
Future Directions. To extend our extreme bandit setting, one could further refine the reward modeling by incorporating the non-stationarity, especially in AutoML for data streams, where optimal models shift with data distributions (Verma et al., 2024). Incorporating cost-aware optimization is another promising direction, as computational resources and time, rather than iteration counts, often define budgets in AutoML; this would require estimating model training times and factoring them into the decision process. Addressing the growing complexity of heterogeneous ML tasks, such as those involving pre-training, fine-tuning, or prompt engineering, may benefit from a hierarchical approach that allocates resources effectively across diverse (Balef and Eggensperger, 2025a, b). Additionally, exploiting structural similarities among algorithms and their hyperparameters could reduce exploration costs by sharing information across arms in the CASH problem, further enhancing efficiency.
Acknowledgments and Disclosure of Funding
The authors are funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC number 2064/1 – Project number 390727645. Additionally, C. Vernade acknowledges funding from the DFG under the project 468806714 of the Emmy Noether Programme. The authors also thank the International Max Planck Research School for Intelligent Systems (IMPRS-IS).
@@ Appendix A Preliminiaries
Let $F(x)=P(X\leq x)$ be the cumulative distribution function, so G(x)=1F(x)=P(X>x)G(x)=1-F(x)=P(X>x). First, consider the probability that the maximum of the $n$ samples is less than or equal to $x$:
@@ Appendix B Proofs
By applying the Mean Value Theorem (MVT) to the survival function $G$ over an interval $b-\epsilon,b]$, there exists a point $c\in(b-\epsilon,b)$ such that:
@@
Where $N_{i}(T)=\mathbbm{E}(\sum\limits_{t=1}^{T}\mathbbm{1}{\{I_{t}=i\}})$ is the number of sub-optimal pulls of arm $i$, arm $i^{*}$ is the optimal arm and $b_{i}$ is the upper bound on the reward of arm $i$ as given by Assumption 3.2, i.e., the reward of arm $i$ lies within the interval $a_{i},b_{i}]$.
@@
$=\left\{\max_{1\leq n\leq n_{1}(t)}r_{1,n}\leq\mathop{\mathbbm{E}[\max_{t\leq T}r_{1,t}]}-\Delta_{i}-C_{t}(n_{1}(t))\right\}$
@@
  • Underestimating the upper confidence bound of extreme values for the optimal arm 11
@@
Now we use Lemma A.1 to calculate the probability of 28:
@@
This selection of the function of the exploration bonus results in two significant advantages. First, it provides an upper bound for the right-hand side of equation 30 that remains independent of $n_{1}$. Furthermore, Equation 31 shows a decreasing trend as the number of pulls for the sub-optimal arm $i$ increases. [^8] We note that this choice is not based on the inherent property of maximum values. In general one can use $C_{t}(n)=(\frac{\alpha\log(t)}{n})^{m}$ for $m>1$ with the optimal $m$ depending on the setting. In Appendix B.4 we show how $m$ affects the regret.
@@
B.4 Extension of Proof of Theorem 4.2
As we discuss in Section 3, the reward distribution in our setting is left-skewed. We now show that under the assumption that the survival function decays rapidly near the maximum i.e., $G_{i}(\mathbbm{E}[\max_{t\leq T}r_{i,t}])=\mathcal{O}(\frac{1}{T^{2}})$, a tighter bound can be derived. This assumption means that the reward distribution has a very light-tail near its upper extreme, which may often hold for our left-skewed distributions. Furthermore, we generalize our algorithm by assuming $C_{t}=(\frac{\alpha\log(t)}{n_{i}})^{m}$ as a exploration bonus function, where $m\geq 1$ is a hyperparameter.
Proof.
We begin with Equation 27 and we take $x=\mathbbm{E}[\max\limits_{t\leq T}r_{i,t}]+c\Delta_{i}$, where $c$ is an arbitrary variable $c\in[0,1]$. We have:
$S$ $\subseteq\left\{\max_{1\leq n\leq n_{1}(t)}r_{1,n}+C_{t}(n_{1}(t))\leq\mathbbm{E}[\max_{t\leq T}r_{i,t}]+c\Delta_{i}\right\}$
$\cup\left\{\max_{1\leq n\leq n_{i}(t)}r_{i,n}+C_{t}(n_{i}(t))>\mathbbm{E}[\max_{t\leq T}r_{i,t}]+c\Delta_{i}\right\}$
$=\left\{\max_{1\leq n\leq n_{1}(t)}r_{1,n}\leq\mathop{\mathbbm{E}[\max_{t\leq T}r_{1,t}]}-(1-c)\Delta_{i}-C_{t}(n_{1}(t))\right\}$
$\cup\left\{\max_{1\leq n\leq n_{i}(t)}r_{i,n}>\mathop{\mathbbm{E}[\max_{t\leq T}r_{i,t}]}-C_{t}(n_{i}(t))+c\Delta_{i}\right\}$
$=\underset{S_{1}}{\underbrace{\left\{\max_{1\leq n\leq n_{1}(t)}r_{1,n}\leq\mathop{\mathbbm{E}[\max_{t\leq T}r_{1,t}]}-(1-c)\Delta_{i}-C_{t}(n_{1}(t))\right\}}}$ (36)
$\cup\underset{S_{2}}{\underbrace{\left\{\max_{1\leq n\leq n_{i}(t)}r_{i,n}>\mathop{\mathbbm{E}[\max_{t\leq T}r_{i,t}]}-C_{t}(n_{i}(t))+c\Delta_{i},\quad C_{t}(n_{i}(t))\leq c\Delta_{i}\right\}}}$ (37)
$\cup\underset{S_{3}}{\underbrace{\left\{\max_{1\leq n\leq n_{i}(t)}r_{i,n}>\mathop{\mathbbm{E}[\max_{t\leq T}r_{i,t}]}-C_{t}(n_{i}(t))+c\Delta_{i},\quad C_{t}(n_{i}(t))>c\Delta_{i}\right\}}}$ (38)
First, we calculate the probability of event $S_{2}$:
$P(S_{2})\leq P(\left\{\max_{1\leq n\leq n_{i}(t)}r_{i,n}>\mathop{\mathbbm{E}[\max_{t\leq T}r_{i,t}]}\right\})\leq n_{i}G_{i}(\mathbbm{E}[\max_{t\leq T}r_{i,t}])\leq TG_{i}(\mathbbm{E}[\max_{t\leq T}r_{i,t}).$ (39)
By calculating the number of sub-optimal pulls of arm $i$ up to time $T$, we know the third part of the event ($S_{3}$) can happen at most $C_{T}^{-1}(c\Delta_{i})=(\frac{\alpha\log(T)}{c\Delta_{i}})^{\frac{1}{m}}$ times:
$N_{i}(T)$ $\leq(\frac{\alpha\log(T)}{c\Delta_{i}})^{\frac{1}{m}}+\sum_{t=1}^{T}P(S_{1})+\sum_{t=1}^{T}P(S_{2})$ (40)
$\leq(\frac{\alpha\log(T)}{c\Delta_{i}})^{\frac{1}{m}}+T^{2}G_{i}(\mathbbm{E}[\max_{t\leq T}r_{i,t}])+\sum_{t=1}^{T}P(S_{1})$ (41)
$\leq(\frac{\alpha\log(T)}{c\Delta_{i}})^{\frac{1}{m}}+M+\sum_{t=1}^{T}P(S_{1})$ (42)
Where $M$ is a constant as we assume $G_{i}(\mathbbm{E}[\max_{t\leq T}r_{i,t}])=\mathcal{O}(\frac{1}{T^{2}})$. Finally, we need to find an upper bound for $P(S_{1})$. We need to differentiate between two situations, when $m=1$ and when $m>1$
For $m=1$. We set $c=1$ and then we have:
$P(S_{1})$ $\leq e^{-n_{1}G_{1}(\mathop{\mathbbm{E}[\max_{t\leq T}r_{1,t}]}-(1-c)\Delta_{i}-C_{t}(n_{1}(t)))}\leq$ (43)
$e^{-L_{1}(C_{t}(n_{1}(t))+(1-c)\Delta_{i})}\leq e^{-\alpha L_{1}\log(T)}\leq T^{-\alpha L_{1}}$ (44)
Finally, we have:
$N_{i}(T)$ $\leq M+\frac{\alpha\log(T)}{\Delta_{i}}+\frac{T^{1-\alpha L_{1}}}{1-\alpha L_{1}}$ (45)
With $\alpha>\frac{1}{L_{1}}$:
$N_{i}(T)=\mathcal{O}(\frac{\log(T)}{L_{1}\Delta_{i}})$ (46)
For $m>1$. We know $n((\frac{a}{n})^{m}+b)\geq ab^{\frac{m-1}{m}}[m(m-1)^{\frac{1}{m}-1}]$. We have:
$P(S_{1})$ $\leq e^{-n_{1}(t)L_{1}(C_{t}(n_{1}(t))+(1-c)\Delta_{i})}\leq e^{-\alpha L_{1}\log(T)((1-c)\Delta_{i})^{\frac{m-1}{m}}[m(m-1)^{\frac{1}{m}-1}]}$ (47)
$\leq T^{-\alpha L_{1}((1-c)\Delta_{i})^{\frac{m-1}{m}}[m(m-1)^{\frac{1}{m}-1}]}$ (48)
For simplicity, we set $c=\frac{1}{2}$. Then:
$N_{i}(T)$ $\leq(\frac{\alpha\log(T)}{c\Delta_{i}})^{\frac{1}{m}}+M+\sum_{t=1}^{T}P(S_{2})$ (49)
$\leq(\frac{2\alpha\log(T)}{\Delta_{i}})^{\frac{1}{m}}+M+\frac{T^{1-\alpha L_{1}(\frac{\Delta_{i}}{2})^{\frac{m-1}{m}}[m(m-1)^{\frac{1}{m}-1}]}}{1-\alpha L_{1}(\frac{\Delta_{i}}{2})^{\frac{m-1}{m}}[m(m-1)^{\frac{1}{m}-1}]}$ (50)
Finally, we have:
$N_{i}(T)=(\frac{2\alpha\log(T)}{\Delta_{i}})^{\frac{1}{m}}+\mathcal{O}(T^{1-\alpha L_{1}(\frac{\Delta_{i}}{2})^{\frac{m-1}{m}}[m(m-1)^{\frac{1}{m}-1}]}).$ (51)
And by choosing $\alpha$ as below
$\alpha=\mathcal{O}\left(\frac{1}{L_{1}(\Delta_{i})^{\frac{m-1}{m}}}\right),$ (52)
We have:
$N_{i}(T)=\mathcal{O}\left(\frac{\log(T)}{L_{1}\Delta_{i}^{{\frac{2m-1}{m}}}}\right)^{\frac{1}{m}}.$ (53)
We would like to emphasize again that this result only holds when $G_{i}(\mathbbm{E}[\max_{t\leq T}r_{i,t}])$ is sufficiently small. By using a weaker assumption, controlling $G_{i}(\mathbbm{E}[\max_{t\leq T}r_{i,t}])$ is necessary. In Theorem B.3, we address this for our method by considering $C_{t}(n_{i}(t))\geq G_{i}(\mathbbm{E}[\max_{t\leq T}r_{i,t}])$ to ensure proper control.
Notably, in general, $\Delta_{i}$ depends on $T$ (see Definition 3.1). Meaning that $\frac{\log(T)}{\Delta_{i}}$ does not necessarily lead to a logarithmic regret. However, in some special scenarios where the reward distributions have different supports, we can ensure that $\Delta_{i}=b_{1}-b_{i}$ as $T$ approaches infinity, and a logarithmic regret is achievable.
Furthermore, Equation 53 shows that the parameter $m$ controls $N_{i}(T)$, the number of times a suboptimal arm is pulled asymptotically. When $T$ is very large, a higher value of $m$ (along with the optimal $\alpha$) improves performance. However, it makes the algorithm more sensitive to the choice of $\alpha$. As shown in Equation 52, with increasing $m$, $\Delta_{i}$ has a greater influence on the optimal $\alpha$. Noting that $\Delta_{i}$ varies across arms and is typically unknown for an unseen task, finding the optimal $\alpha$ is not feasible in practice. Therefore, there is a trade-off between performance and sensitivity. In the CASH setting, $m=2$ performs empirically well, while exhibiting low sensitivity to $\alpha$. Furthermore, as shown in Appendix D.4, we can find a range of $\alpha$ for which the MaxUCB works well across different CASH tasks.